Greenhab Report – April 22nd

Crew 245 GreenHab Report 22-04-2021

GreenHab Officer: none, report written by Crew Scientist Olivia Ettlin

Environmental control: heater is kept on all day at 65 degrees.

Average temperatures: heat kept to 65 degrees, daily temperatures higher due to sunlight.

Hours of supplemental light: 0 hours

Daily water usage for crops: none

Daily water usage for research and/or other purposes: add 500 mL to hydroponics garden daily, 1000mL added to regolith feasibility experiment

Water in Blue Tank: 0 gallons

Time(s) of watering for crops: regolith experiment crops watered at approximately 11:45 this morning. It only took half a watering can (so approx. 1000mL) to re-hydrate the seeds.

Changes to crops: no crops to maintain

Narrative: Today, we noticed that the cotyledons that were emerging from the soil in the regolith feasibility experiment were starting to dry up. The seeds themselves still have viable roots, but this is more proof to us that the extreme heat of the season might not have been conducive to the growth of young plants. This wasn’t really apparent while the seeds were germinating, but we’ll know for next time to leave young plants in the lower deck of the Hab, and save the Greenhab for our established plants. We’ll also make sure for next time to monitor the temperature better during the day and implement more environmental controls as needed.

In exciting news, the hydroponics experiment has been a resounding success! The garden yielded 20 tall green onion plants and 16 tomato plants with hearty roots. I’ll soon start crunching the numbers to determine if we had any significant difference in plant growth between the fertilizer variables, but overall this experiment has provided strong evidence to support the feasibility of hydroponics gardens on Mars.

Harvest: none

Support/supplies needed: none at this time.

Operations Report – April 22nd

Crew 245 Operations Report 22-04-2021

SOL: 12

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Working nominally

Hours run: 7.5

From what time last night: 9:00 pm, SOC 60%

To what time this morning: 4:30 am, SOC 100%

List any additional daytime hours when the generator was run: N/A

Solar— SOC 73% at 7:28 pm (Before generator is run at night)

Notes on power system: For the last several sols, the freezers in the Science Dome have been running in order to preserve science samples. Due to this, the station power consumption has been significantly higher than it previously was, resulting in the generator being turned on earlier (as early as 7:00 pm on colder days, and 9:00 pm on warmer days) once the SOC reaches below 70%.

Diesel Reading – 50%

Station Propane Reading – 70%

Water (loft tank): 40 gallons

Water Meter: 150725.5 units

Water (static tank): 65 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: No

Hours: 204.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Left tail-light on Perseverance rover is not functioning – noticed on EVA.

ATV’s Used: None

Reason for use: N/A

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were used today: 0

Notes on ATVs: Mission Director will monitor and refuel ATVs for the duration of the mission.

HabCar used and why, where?: Nothing to report

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Internet is working nominally. Crew is rationing internet usage during the day to ensure that there is sufficient bandwidth for Mission Support Comms window.

Summary of suits and radios: All radios nominal, all suits currently charging. Suits 7, 9, and 11 shifted to different outlet, as initial power strip was not charging suits properly. Suits 3, 4, and 6 used on EVA today, and all functioned nominally.

Summary of Hab operations: Nothing to report besides nominal crew activities.

Summary of GreenHab operations: Being used to host Crew Botanist botany experiment, but no active operations.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing to report

Shravan Hariharan, Crew Engineer
Red Planet People – MDRS Crew 245 "Team Patamars"
To Mars and Beyond – For All!

Sol Summary – April 22nd

Crew 245 Sol Summary Report 22-04-2021

Sol: 12

Summary Title: Final Steps Outside the Hab

Author’s name: Shayna Hume

Mission Status: Executive Officer

Sol Activity Summary:

· 08:30-09:30: Misc. breakfast and briefing

· 09:30-10:00: Prepare for EVA.

· 10:00-12:00: EVA.

· 12:00-13:30: Lunch!

· 13:30-18:00: Misc. tasks. Videos were taken of crew members explaining their experiments. Interviews were done for the post-mission mockumentary. Final soil samples were analyzed. Parts of the Hab were cleaned in preparation for checkout at the end of the mission.

· 18:00-19:00: Dinner

Look Ahead Plan: Tomorrow we intend to spend the morning wrapping up a majority of our research tasks, while in the afternoon we intend to record some media for our mockumentary and spend several hours cleaning in preparation for the mission checkout.

Anomalies in work: None.

Weather: Nominal.

Crew Physical Status: No issues.

EVA: EVA #11 report coming in.

Reports to be filed: Operations Report, EVA Report, GreenHab Report

Support Requested: n/a

Shayna Hume, Executive Officer
Red Planet People – MDRS Crew 245 "Team Patamars"
To Mars and Beyond – For All!

EVA Report – April 22nd

Crew 245 EVA Report 04-22-2021

EVA # 11

Author: Shayna Hume, Executive Officer

Purpose of EVA: Gathering soil samples on a ridge area.

Start time: 1000 MDT

End time: 1200 MDT (planned 1300, actual 1200)

Narrative: The EVA this morning started on schedule after the crew had carefully gathering and sterilizing required scientific instruments in order to meet the crew’s determined Planetary Protection Protocols. After suiting up, the crew waited 5 minutes before exiting. Today, the crew departed from the Hab and Hab Grounds by foot, heading northwest.

After a thirty-minute hike, we reached the top of the western ridge and began heading North towards Schubert’s Pass. After taking soil samples along the ridge, the team took a few minutes to admire the incredible view of North Ridge, Schubert’s Pass, and the western buttes.

After determining there were no valid candidates for emergency shelter, the team headed back along the 1103 route pointed out by the outpost as advice. The team managed a very quick return by foot to the Hab and was in the airlock just before 1200.

Destination: Schubert’s Pass

Participants: Executive Officer, Botanist, Engineer

Road(s) and routes per MDRS Map: Walking up near a ridge slope by the Hab, taking 1103 walking back.

Mode of travel: Walking.
Shayna Hume, Executive Officer
Red Planet People – MDRS Crew 245 "Team Patamars"
To Mars and Beyond – For All!

Copyright © The Mars Society. All rights reserved. | Main Site